OpenStudy (anonymous):

Snookers lumber can convert logs into either lumber or plywood. In a given day, the mill turns out twice as many units of plywood as lumber. It makes a profit of $30 on a unit of lumber and $50 on a unit of plywood. How many of each unit must be produced and sold in order to make a profit of $13,130

OpenStudy (anonymous):

let the units of lumber = x then no of units of plywood = 2x profit on x units of lumber = x*30 = 30x profit on 2x units of plywood = 2x*50 = 100x ATQ, 30x + 100x = 13,130 Now solve for x ..... n get units of lumber and 2x= no of units of plywood

OpenStudy (anonymous):

say if u need help

OpenStudy (anonymous):

what is the ATQ?

OpenStudy (anonymous):

I have to log off for a bit. I will come back to this one later.